### v2

# On kinks and solitons in discrete nonlinear transmission line

Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons. We calculated the speeds and the profiles both of the kinks and of the solitons. We also studied dissipative kinks in the lossy transmission lines, with ohmic resistors shunting the inductors. The losses influence the dissipative kink speed and profile.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . Among the latter – the travelling waves, especially the kinks and the solitons, attract a lot of attention solitons . The nonlinear waves become especially interesting (and challenging) when one takes into account the discrete nature of the transmission line. The kinks and the solitons in discrete transmission lines were studied in quite a few publications malomed4 ; cameroon ; bogdan .
The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi . In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein
We studied previously the kinks and the solitons in discrete Josephson transmission line (JTL) JTL kogan2 ; kogan3 . There we formulated the quasi-static approximation, which allows to reduce the discrete equations to the continuous ones. In the present short note we would like to apply this approximation to different kind of discrete transmission lines – with nonlinear capacitors. Such transmission lines in the continuum approximation were studied in the classical papers by R. Landauer landauer1 ; landauer2 ; landauer .
II The kinks and the solitons
II.1 The travelling waves
The discrete transmission line is constructed from the identical inductors and nonlinear capacitors is shown on Fig. 1.
Figure 1: Discrete transmission line.
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations are
$\frac{dQ_{n}}{dt}$ $=I^{c}_{n},$ (1a)
$L\frac{dI^{c}_{n}}{dt}$ $=V_{n+1}-2V_{n}+V_{n-1},$ (1b)
where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the function $Q(V)$. We also took into account the connection between the current through the capacitor $dI^{c}_{n}$ and the currents between the inductors
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}.$ (2)
Let us write down Eq. (II.1) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 . The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s. of the equation (1b) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
After that (II.1) takes the form
$\frac{\partial Q}{\partial t}$ $=I^{c},$ (4a)
$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}\right).$ (4b)
In this paper we are interested in the travelling wave solutions of (II.1), that is the dependence of all the variables upon $Z$ and $t$ is of the form
$X(Z,t)=X(Z-Ut),$ (5)
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This gives us the opportunity to perform integration in (4b) to get
$ULI^{c}=\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
Substituting (6) into (4a) we obtain ordinary differential equation
$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
An obvious way to integrate (7) starts by going to the second order ODE
$LU^{2}Q=V+\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\text{const}.$ (8)
However in this paper we’ll go along the different way. Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
$E=\frac{dV}{dZ}$ (9)
and considering $V$ as the independent variable, we may after taking into account that
$\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (10a)
$\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (10b)
and dividing both parts by the common multiplier $E$ in both parts, reduce the order of (7) and write it down as
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\frac{dQ(V)}{dV}-1.$ (11)
A simple mechanical interpretation of Eq. (11) may be helpful. If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s. of (11).
We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
$\lim_{Z\to-\infty}V(Z)=V_{1},\hskip 28.45274pt\lim_{Z\to+\infty}V(Z)=V_{2}.$ (12)
The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton.
The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (11) follows
$LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (13)
The inequality (13) has a simple physical meaning. Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$. These waves are described by the equation obtained by linearizing (7) near $V=V_{1,2}$ and ignoring the term with the third order derivative
$Lu^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}\frac{dV}{dZ}=\frac{dV}{dZ}.$ (14)
Hence the speed of the sound is
$\frac{1}{Lu_{1,2}^{2}}=\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}.$ (15)
Now we understand that (13) means
$U^{2}>u_{1,2}^{2},$ (16)
that is the travelling waves are supersonic kogan2 ; kogan3 .
The boundary conditions (12) obviously give
$E\left(V_{1,2}\right)=0.$ (17)
From (10a) follows
$\left(\frac{dE^{2}}{dV}\right)_{V=V_{1,2}}=0.$ (18)
The conditions (17) and (18) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
In the vicinity of $V_{1,2}$ Eq. (11) has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (19)
Taking into account the boundary conditions (17) and (18) we obtain
$E^{2}\sim\left(V-V_{1,2}\right)^{2},\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (20)
If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (21)
Eq. (11) can be specified to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (22)
For brevity we’ll call Eq. (21) and (22) the cubic case.
II.2 The kinks
With four boundary conditions on $E(V)$ supplied by (17) and (18) for the kink problem of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two being connected with it. One such connection we could obtain at much earlier stage. In fact, integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
$\frac{1}{LU^{2}}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (23)
where $Q_{i}=Q(V_{i})$.
If $Q(V)$ is a polynomial function of $V$, from (20) follows that the kink can be presented as
$E^{2}(V)=P_{1}(V)(V_{1}-V)^{2}(V-V_{2})^{2}.$ (24)
In the cubic case, calculating the integral in (23) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right).$ (25)
We also can specify (24) to
$E^{2}(V)=\psi^{2}(V_{1}-V)^{2}(V-V_{2})^{2},$ (26)
where $\psi$ is a constant. Substituting (26) into (22) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (27a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}.$ (27b)
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (28)
The solution is
$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right]$ (29)
(to satisfy the boundary conditions we have to choose positive value of $\psi$). After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2LU}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (30)
II.3 The solitons
To specify the soliton we need to introduce an additional parameter, and as such we’ll take the extremal value of the voltage $V_{3}$. That is $V$ changes between $V_{1}$ and $V_{3}$, and $|V_{3}-V_{1}|$ we’ll call the amplitude of the soliton. Because $V_{3}$ is the extremal value of $V$ we have
$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0\Rightarrow E(V_{3})=0.$ (31)
Hence
$E^{2}\sim\left(V-V_{3}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{3}.$ (32)
Recalling our mechanical analogy we see that the soliton corresponds to the motion of the fictitious particle with zero initial coordinate and velocity at $V=V_{1}$ and zero coordinate (but not velocity) at $V=V_{3}$. Mechanical interpretation of (11) in the case of the soliton leads to the situation which doesn’t normally occur in mechanics. The boundary condition $V_{2}=V_{1}$ means that the ”time” $V$ flows along the contour from $V_{1}$ to $V_{3}$ and then back from $V_{3}$ to $V_{1}$. This is why $E(V)$ for the soliton will later turn out to be a double-valued function.
With three boundary conditions on $E(V)$ supplied by (17), (18) and (31) for the soliton problem of the three parameters $V_{1}V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
[…]
If $Q(V)$ is a polynomial function of $V$, from (20) and (32) follows that the soliton can be presented as
$E^{2}(V)=P_{2}(V)(V_{1}-V)^{2}(V-V_{3}),$ (34)
where $P_{1}(V)$ and $P_{2}(V)$ are some polynomial functions.
In the cubic case, calculating the integral in (33) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right).$
We also can specify (34) (generalize (26)) to
$E^{2}(V)=\psi(V_{1}-V)^{2}(V-V_{3})(V-V_{4}),$ (36)
where $\psi$ is a constant. Substituting (26) into (22) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}=6U^{2}LC_{0}\gamma,$ (37a)
$2V_{1}+V_{3}+V_{4}=-\frac{4\beta}{3\gamma}.$ (37b)
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (21). In this case, excluding $V_{4}$, we obtain that (36) and (II.3) are modified to
$E^{2}(V)$ $=\phi(V_{1}-V)^{2}(V-V_{3}),$ (38a)
$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta,$ (38b)
$\frac{1}{LC_{0}U^{2}}$ $=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right).$ (38c)
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$). From (38c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation
$\frac{1}{LC_{0}u^{2}}=\alpha+2\beta V_{1},$ (39)
in line with our previous general statement that the soliton is supersonic.
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (40)
The solution of (40) is
$V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (41)
Similar to what was done in the previous subsection, we can use (1b) to find the current and obtain
$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2LU}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (42)
II.4 The quasi-continuum approximation
We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper. To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s. of (1b) kogan2 ; kogan3
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}+\frac{1}{360}\frac{d^{6}V}{dZ^{6}}.$ (43)
Thus we obtain instead of (7) the equation
$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (44)
Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
The small numerical coefficient before the sixth order derivative in the r.h.s. of (44) means that the small parameters shouldn’t be actually very small. With all that, one can expect that Eq. (44) will allow to study the cases of stronger nonlinearity than Eq. (7).
III The dissipative kinks
Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
Figure 2: Discrete transmission line with shunted inductors.
In this case (2) is modified to
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}+\frac{1}{R}\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (45)
hence (II.1) takes the form
$\frac{\partial Q}{\partial t}$ $=I^{c},$ (46a)
$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}+\frac{L}{R}\frac{\partial^{2}V}{\partial t\partial Z}\right).$ (46b)
(we assume that both the higher order derivatives with respect to $Z$ and $1/R$ are small, so the discrete second order derivative with respect to $Z$ in the r.h.s. of (45) we presented as the continuous second order derivative).
After repeating exactly what was done in Section II.1 we obtain in a modified form Eq. (11)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}\frac{dQ(V)}{dV}-1$ (47)
and (in the cubic case) Eq. (22)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (48)
We again impose on $V(Z)$ the boundary conditions (12), which lead to the boundary conditions (17) and (18) for $E(V)$. Equation (47) in the vicinity of $V_{1,2}$ has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=\text{const},$ (49)
which contains additional term in the l.h.s. with respect to (19), but we still obtain the condition (20). We can also repeat verbatim the part of Section II.2 leading to Eqs. (23) and (24).
In the cubic case we regain Eq. (25), and there exists kogan4 a simple polynomial solution of (48) in the form (26), which we repeat here
$E(V)=\psi(V_{1}-V)(V-V_{2}),$ (50)
where $\psi$ is a constant. Substituting (50) into (48) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (51a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4LU}{\psi R}..$ (51b)
Note that Eqs. (27a), (37a) and (51a) are all identical, hence all the waves we consider in this paper can exist only if $\gamma>0$. Note also that by integrating (48) we can obtain Abel equation of the second kind, and (51b) is the condition of the solvability of the equation polyanin .
Equation (50) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (29)). However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (51b) instead of (27b)) and so does the kink speed, in spite of the validity in both cases of Eq. (25).
There exists different from (50) type of solutions of (47) (hence different from (29) type of solutions of (III), similar to the shock waves we considered previously kogan1 ; kogan2 ; kogan3 ; kogan4 . In the present paper we don’t consider such solutions.
IV Conclusions
We considered finite support travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied the dissipative kinks in the lossy discrete transmission line. The losses influence the kink speed and profile.
Acknowledgements.
I am grateful to B. Malomed for the discussion. The paper was finalized during the Author’ participation in Workshop on Twistronics and Moire Materials: Bridging Theory and Experiments, ICTP, Trieste. The Author is grateful to the Center for the hospitality.

### v3

@@
-On kinks and solitons in discrete nonlinear transmission line
+On the kinks, the solitons and the shocks in discrete nonlinear transmission line
 Abstract
 We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors.
-The travelling waves with finite support in such systems are the kinks and the solitons.
-We calculated the speeds and the profiles both of the kinks and of the solitons.
-We also studied dissipative kinks in the lossy transmission lines, with ohmic resistors shunting the inductors.
-The losses influence the dissipative kink speed and profile.
+The travelling waves with finite support in such systems are the kinks and the solitons, which speeds and profiles were calculated.
+We also studied lossy transmission lines.
+The travelling waves with finite support in such systems are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
 I Introduction
@@
 Figure 1: Discrete transmission line.
-We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations are
+We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
 $\frac{dQ_{n}}{dt}$ $=I^{c}_{n},$ (1a)
-$L\frac{dI^{c}_{n}}{dt}$ $=V_{n+1}-2V_{n}+V_{n-1},$ (1b)
-where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the function $Q(V)$.
-We also took into account the connection between the current through the capacitor $dI^{c}_{n}$ and the currents between the inductors
-$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}.$ (2)
+$L\frac{dI_{n,n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
+$I^{c}_{n}$ $=I_{n-1,n}-I_{n,n+1}.$ (1c)
+where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the (monotonically increasing) function $Q=Q(V)$.
+We can exclude $I_{n,n+1}$ from (1b) and (1c) and present these two equations a single one
+$L\frac{dI^{c}_{n}}{dt}=V_{n+1}-2V_{n}+V_{n-1}.$ (2)
 Let us write down Eq.
-(II.1) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 .
+(2) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 .
 The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s.
-of the equation (1b) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion
-$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
-After that (II.1) takes the form
-$\frac{\partial Q}{\partial t}$ $=I^{c},$ (4a)
-$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}\right).$ (4b)
-In this paper we are interested in the travelling wave solutions of (II.1), that is the dependence of all the variables upon $Z$ and $t$ is of the form
+of the equation (2) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion, thus obtaining the equation in the form
+$L\frac{\partial I^{c}_{n}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
+(And, of course, in the framework of the approximation the derivative in (1a) should be understood as partial derivative.)
+From (1a) and (3) we can obtain closed wave equation for the voltage $V$
+$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (4)
+However, further on we will not use Eq. (4); it will be more convenient to use Eqs. (1a) and (3).
+In this paper we are interested in the travelling wave solutions, for which the dependence of all the variables upon $Z$ and $t$ is of the form
 $X(Z,t)=X(Z-Ut),$ (5)
 where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$).
-This gives us the opportunity to perform integration in (4b) to get
-$ULI^{c}=\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
+This gives us the opportunity to perform integration in (3) to get
+$ULI^{c}=-\frac{\partial V}{\partial Z}-\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
 Substituting (6) into (1a) we obtain ordinary differential equation
-$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
-An obvious way to integrate (7) starts by going to the second order ODE
-$LU^{2}Q=V+\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\text{const}.$ (8)
-However in this paper we’ll go along the different way. Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
+$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
+Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
 $E=\frac{dV}{dZ}$ (8)
@@
 and dividing both parts by the common multiplier $E$ in both parts, reduce the order of (7) and write it down as
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\frac{dQ(V)}{dV}-1.$ (11)
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\frac{dQ(V)}{dV}-1.$ (10)
 A simple mechanical interpretation of Eq.
@@
 The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (10) follows
-$LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (13)
+$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (12)
 The inequality (12) has a simple physical meaning.
 Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$.
-These waves are described by the equation obtained by linearizing (7) near $V=V_{1,2}$ and ignoring the term with the third order derivative
-$Lu^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}\frac{dV}{dZ}=\frac{dV}{dZ}.$ (14)
-Hence the speed of the sound is
-$\frac{1}{Lu_{1,2}^{2}}=\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}.$ (15)
+These waves are described by Eq.
+(10), with the l.h.s.
+being put equal to zero.
+Hence the speed of the sound is given by the equation
+$\frac{1}{u^{2}L}=\frac{dQ}{dV}.$ (13)
 Now we understand that (12) means
-$U^{2}>u_{1,2}^{2},$ (16)
+$U^{2}>u^{2}\left(V_{1,2}\right),$ (14)
 that is the travelling waves are supersonic kogan2 ; kogan3 .
@@
 $E\left(V_{1,2}\right)=0.$ (15)
-From (10a) follows
-$\left(\frac{dE^{2}}{dV}\right)_{V=V_{1,2}}=0.$ (18)
-The conditions (17) and (18) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
+The conditions (15) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
 In the vicinity of $V_{1,2}$ Eq. (10) has the form
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (19)
-Taking into account the boundary conditions (17) and (18) we obtain
-$E^{2}\sim\left(V-V_{1,2}\right)^{2},\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (20)
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (16)
+Taking into account the boundary conditions (15) we obtain
+$E\sim\left(V-V_{1,2}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (17)
 If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
@@
 Eq. (10) can be specified to
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (22)
-For brevity we’ll call Eq. (21) and (22) the cubic case.
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (19)
+For brevity we’ll call Eq. (18) and (19) the cubic case. Not that in this case
+$\frac{1}{LC_{0}u^{2}(V)}=\alpha+2\beta V+3\gamma V^{2}.$ (20)
 II.2 The kinks
@@
 In fact, integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
-$\frac{1}{LU^{2}}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (23)
+$\frac{1}{U^{2}L}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (21)
 where $Q_{i}=Q(V_{i})$.
 If $Q(V)$ is a polynomial function of $V$, from (17) follows that the kink can be presented as
-$E^{2}(V)=P_{1}(V)(V_{1}-V)^{2}(V-V_{2})^{2}.$ (24)
+$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2}.$ (22)
 In the cubic case, calculating the integral in (21) we obtain
@@
 We also can specify (22) to
-$E^{2}(V)=\psi^{2}(V_{1}-V)^{2}(V-V_{2})^{2},$ (26)
+$E^{2}(V)=\psi^{2}(V-V_{1})^{2}(V-V_{2})^{2},$ (24)
 where $\psi$ is a constant. Substituting (24) into (19) and using Vieta’s formula we obtain after simple algebra
@@
 The solution is
-$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right]$ (29)
-(to satisfy the boundary conditions we have to choose positive value of $\psi$). After finding $V(Z)$ we can from (1b) obtain the current
-$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2LU}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (30)
+$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (27)
+After finding $V(Z)$ we can from (2) obtain the current
+$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2UL}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (28)
 II.3 The solitons
@@
 This is why $E(V)$ for the soliton will later turn out to be a double-valued function.
-With three boundary conditions on $E(V)$ supplied by (17), (18) and (31) for the soliton problem of the three parameters $V_{1}V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters.
+With three boundary conditions on $E(V)$ supplied by (15) and (29) for the soliton problem of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters.
 Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
@@
 If $Q(V)$ is a polynomial function of $V$, from (17) and (30) follows that the soliton can be presented as
-$E^{2}(V)=P_{2}(V)(V_{1}-V)^{2}(V-V_{3}),$ (34)
+$E^{2}(V)=P_{2}(V)(V-V_{1})^{2}(V-V_{3}),$ (32)
 where $P_{1}(V)$ and $P_{2}(V)$ are some polynomial functions.
@@
 We also can specify (32) (generalize (24)) to
-$E^{2}(V)=\psi(V_{1}-V)^{2}(V-V_{3})(V-V_{4}),$ (36)
+$E^{2}(V)=\psi(V-V_{1})^{2}(V-V_{3})(V-V_{4}),$ (34)
 where $\psi$ is a constant. Substituting (24) into (19) and using Vieta’s formula we obtain after simple algebra
@@
 In this case, excluding $V_{4}$, we obtain that (34) and (II.3) are modified to
-$E^{2}(V)$ $=\phi(V_{1}-V)^{2}(V-V_{3}),$ (38a)
+$E^{2}(V)$ $=\phi(V-V_{1})^{2}(V-V_{3}),$ (36a)
 $\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta,$ (36b)
@@
 For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
-From (38c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation
-$\frac{1}{LC_{0}u^{2}}=\alpha+2\beta V_{1},$ (39)
-in line with our previous general statement that the soliton is supersonic.
+From (36c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation (20) with $\gamma=0$, in line with our previous general statement that the soliton is supersonic.
 To find $V(Z)$ we have to solve equation
@@
 $V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (38)
-Similar to what was done in the previous subsection, we can use (1b) to find the current and obtain
-$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2LU}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (42)
+Similar to what was done in the previous subsection, we can use (2) to find the current and obtain
+$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (39)
 II.4 The quasi-continuum approximation
@@
 To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s.
-of (1b) kogan2 ; kogan3
+of (2) kogan2 ; kogan3
 $V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}+\frac{1}{360}\frac{d^{6}V}{dZ^{6}}.$ (40)
 Thus we obtain instead of (7) the equation
-$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (44)
+$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (41)
 Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
@@
 (7).
-III The dissipative kinks
+III The dissipative kinks and shocks
 Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
 Figure 2: Discrete transmission line with shunted inductors.
-In this case (2) is modified to
+In this case (1c) is modified to
 $I^{c}_{n}=I_{n-1,n}-I_{n,n+1}+\frac{1}{R}\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (42)
-hence (II.1) takes the form
-$\frac{\partial Q}{\partial t}$ $=I^{c},$ (46a)
-$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}+\frac{L}{R}\frac{\partial^{2}V}{\partial t\partial Z}\right).$ (46b)
+hence Eq. (3) takes the form
+$L\frac{\partial I^{c}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}-\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial Z^{2}}.$ (43)
 (we assume that both the higher order derivatives with respect to $Z$ and $1/R$ are small, so the discrete second order derivative with respect to $Z$ in the r.h.s.
 of (42) we presented as the continuous second order derivative).
-After repeating exactly what was done in Section II.1 we obtain in a modified form Eq. (11)
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}\frac{dQ(V)}{dV}-1$ (47)
-and (in the cubic case) Eq. (22)
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (48)
-We again impose on $V(Z)$ the boundary conditions (12), which lead to the boundary conditions (17) and (18) for $E(V)$. Equation (47) in the vicinity of $V_{1,2}$ has the form
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=\text{const},$ (49)
+From (1a) and (43) we can obtain Eq. (4) modified due to the presence of ohmic resistors
+$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (44)
+For the travelling wave, combining (1a) and (43), we obtain after integration closed equation for $V(Z)$
+$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=U^{2}LQ-V+F,$ (45)
+where $F$ is the constant of integration.
+Equation (45) is Newton equation, describing motion with friction of the fictitious particle ($V$ being the coordinate of the particle and $Z$ the ”time”) in the force field given by the r.h.s.
+of the equation.
+We again impose on $V(Z)$ the boundary conditions (11).
+Hence the motion starts at $Z=-\infty$ at $V=V_{1}$ and ends at $Z=+\infty$ at $V=V_{2}$ (obviously that both $V_{1}$ and $V_{2}$ are the equilibrium positions).
+Equating the r.h.s.
+of (45) to zero at $V=V_{1}$ and $V=V_{2}$ we recover Eq.
+(21) for the speed of the wave.
+It will be convenient to present Eq. (45) as
+$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (46)
+where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
+$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (47)
+The constraint for the travelling wave asymptotic boundary conditions is the inequality
+$\Pi(V_{1})>\Pi(V_{2}).$ (48)
+Inequality (48) can be explicitly presented as
+[…]
+When both $\beta>0$ and $\gamma>0$, the condition (49) means just $|V_{1}|>|V_{2}|$.
+The point $V=V_{1}$ should be the position of unstable equilibrium.
+The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium.
+We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ).
+The stable equilibrium waves we called shocks, and the unstable equilibrium - kinks.
+There are several important distinctions between these two types of waves.
+Unstable equilibrium at both at the point $V=V_{1}$ and $V=V_{2}$ means that differentiating the r.h.s.
+of (45) we recover equation (12).
+The kink, similar to what was in the lossless case, is supersonic both with respect to the region before the kink and the region behind the kink.
+Since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation.
+Stable equilibrium at the point $V=V_{2}$ means
+$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{2}}-1<0.$ (50)
+Inequalities (50) and (12) (the latter only for $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
+To clarify the distinction between the shocks and the kinks let us bring together the formulas for the travelling wave speed and the sound speed for the cubic case
+$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right),$ (51a)
+$\frac{1}{LC_{0}u^{2}(V_{1,2})}=\alpha+2\beta V_{1,2}+3\gamma V_{1,2}^{2}.$ (51b)
+We realize that for $\beta>0$ and $\gamma>0$, for $V_{1}$ and $V_{2}$ having the same sign the travelling wave is always the shock.
+For $V_{1}$ and $V_{2}$ having opposite signs, we can have both the kink and the shock, depending upon the parameters.
+To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (10)
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L\frac{dQ(V)}{dV}-1.$ (52)
+The boundary conditions (11) lead to the boundary conditions (15) for $E(V)$, like it was in the lossless case. Equation (52) in the vicinity of $V_{1,2}$ has the form
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=\text{const},$ (53)
 which contains additional term in the l.h.s. with respect to (16), but we still obtain the condition (17). We can also repeat verbatim the part of Section II.2 leading to Eqs. (21) and (22).
-In the cubic case we regain Eq. (25), and there exists kogan4 a simple polynomial solution of (48) in the form (26), which we repeat here
-$E(V)=\psi(V_{1}-V)(V-V_{2}),$ (50)
-where $\psi$ is a constant. Substituting (50) into (48) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
+In the cubic case Eq. (52) becomes
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (54)
+We regain Eq. (23), and there exists kogan4 a simple polynomial solution of (54) in the form (24), which we repeat here
+$E(V)=\psi(V-V_{1})(V-V_{2}),$ (55)
+where $\psi$ is a constant (positive for positive $V_{1}$).
+Substituting (55) into (54) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
 $\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (56a)
-$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4LU}{\psi R}..$ (51b)
+$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4UL}{\psi R}.$ (56b)
 Note that Eqs.
@@
 Equation (55) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (27)).
+And the the shocks are of the same shape also.
 However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq.
-(51b) instead of (27b)) and so does the kink speed, in spite of the validity in both cases of Eq.
+(56b) instead of (25b)) and so does the kink (the shock) speed, in spite of the validity in all the cases of Eq.
 (23).
-There exists different from (50) type of solutions of (47) (hence different from (29) type of solutions of (III), similar to the shock waves we considered previously kogan1 ; kogan2 ; kogan3 ; kogan4 .
-In the present paper we don’t consider such solutions.
+The boundary in the phase space between the kinks and the shocks can be easily obtained from (III) in the case when $\beta=0$. Equating the speeds we obtain that this boundary corresponds to
+$V_{1}+2V_{2}=0$ (57)
+(very much similar to what we had for the Josephson transmission line kogan4 ).
+We need to emphasise the different meaning of the condition (56b) for the kinks and the shocks.
+For the kinks, (56b) can be understood as the constraint on $V_{2}$ (for a given value of $V_{1}$ and ohmic resistance $R$).
+And it happens that the kink satisfying such constraint is described by the elementary function.
+For the shock, (56b) is just the condition for it being described by an elementary function.
+But, of course, the shock with any other value of $V_{2}$ (satisfying (48) also exists.
+Let us assume that $V_{1}$ and $V_{2}$ are so close to each other, that we can keep in the r.h.s. of (18) only the first two terms, thus simplifying (54) to
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1.$ (58)
+The solution of (58) satisfying the boundary conditions (15) and (29) is
+$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right],$ (59)
+where $\Delta V\equiv V_{1}-V_{2}$ (we assumed for the sake of definiteness that $V_{1},V_{2}>0$). Substituting (59) into (58) we obtain
+$\chi^{2}$ $=8U^{2}LC_{0}\beta$ (60a)
+$\frac{UL}{R}$ $=\frac{5\chi}{24}\left(\Delta V\right)^{1/2}.$ (60b)
+The function $E(V)$ being found, we can find the function $V(Z)$ by solving equation
+$\frac{dV}{dX}=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right].$ (61)
+The solution is
+$V(Z)=V_{2}+\frac{\Delta V}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (62)
+According to what was said after Eq.
+(III), Eq.
+(62) describes the shock.
+And we have to repeat here what was said after Eq.
+(III).
+Equation (60b) is just the condition for the shock to be expressed by an elementary function.
+The shock exists for any other value of $V_{2}$.
 IV Conclusions
@@
 We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line.
-We also studied the dissipative kinks in the lossy discrete transmission line.
-The losses influence the kink speed and profile.
+We also studied lossy transmission lines.
+The travelling waves with finite support in such systems are the dissipative kinks and the shocks.
+We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.
 Acknowledgements.

### v5

@@
 The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi .
-In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein
+In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein ; vainchtein2 ; kamchatnov
 We studied previously the kinks and the solitons in discrete Josephson transmission line (JTL) JTL kogan2 ; kogan3 .
@@
 $L\frac{dI^{c}_{n}}{dt}=V_{n+1}-2V_{n}+V_{n-1}.$ (2)
+From (1a) and (2) we can obtain closed wave equation for the voltage $V$
+$L\frac{d^{2}Q_{n}}{dt^{2}}=V_{n+1}-2V_{n}+V_{n-1}.$ (3)
 Let us write down Eq.
@@
 of the equation (3) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion, thus obtaining the equation in the form
-$L\frac{\partial I^{c}_{n}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
-(And, of course, in the framework of the approximation the derivative in (1a) should be understood as partial derivative.)
-From (1a) and (3) we can obtain closed wave equation for the voltage $V$
 $L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (4)
-However, further on we will not use Eq. (4); it will be more convenient to use Eqs. (1a) and (3).
 In this paper we are interested in the travelling wave solutions, for which the dependence of all the variables upon $Z$ and $t$ is of the form
 $X(Z,t)=X(Z-Ut),$ (5)
-where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$).
-This gives us the opportunity to perform integration in (3) to get
-$ULI^{c}=-\frac{\partial V}{\partial Z}-\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
-Substituting (6) into (1a) we obtain ordinary differential equation
+where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This from (4) we obtain an ordinary differential equation
+$U^{2}L\frac{d^{2}Q(V)}{dZ^{2}}=\frac{d^{2}V}{dZ^{2}}+\frac{1}{12}\frac{d^{4}V}{dZ^{4}}.$ (6)
+Performing an integration we get
 $U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
-Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
-$E=\frac{dV}{dZ}$ (8)
-and considering $V$ as the independent variable, we may after taking into account that
+One way to solve Eq. (7) is to perform the second integration to obtain
+$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}=U^{2}LQ(V)-V+F,$ (8)
+where $F$ is the constant of integration.
+However, in this Section we’ll use another method of solution of Eq. (7). After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, we are motivated to introduce the new dependent variable
+$E=\frac{dV}{dZ}.$ (9)
+Considering $V$ as the independent variable, taking into account that
 $\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (10a)
 $\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (10b)
-and dividing both parts by the common multiplier $E$ in both parts, reduce the order of (7) and write it down as
+and dividing both parts by the common multiplier $E$ in both parts, we may reduce the order of (7) and write it down as
 $\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\frac{dQ(V)}{dV}-1.$ (11)
+Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V$.
+These waves are described by Eq.
+(11), with the l.h.s.
+being put equal to zero.
+The speed of the sound is given by the equation
+$\frac{1}{u^{2}(V)}=L\frac{dQ}{dV}.$ (12)
+Thus Eq. (11) can be presented as
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=\frac{U^{2}}{u^{2}(V)}-1.$ (13)
+Note that in the framework of the quasi-continuum approximation $E$ is just the voltage on the inductor.
+Thus the physical meaning of Eq.
+(13) is simple.
+Its solution gives the voltage on the inductor as the function of the voltage on the capacitor.
 A simple mechanical interpretation of Eq.
@@
 If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s.
-of (10).
+of the equation.
 We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
@@
 The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton.
-The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (10) follows
-$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (12)
-The inequality (12) has a simple physical meaning.
-Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$.
-These waves are described by Eq.
-(10), with the l.h.s.
-being put equal to zero.
-Hence the speed of the sound is given by the equation
-$\frac{1}{u^{2}L}=\frac{dQ}{dV}.$ (13)
-Now we understand that (12) means
+The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$.
+Hence from (13) follows
 $U^{2}>u^{2}\left(V_{1,2}\right),$ (15)
@@
 $E\left(V_{1,2}\right)=0.$ (16)
-The conditions (15) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
-In the vicinity of $V_{1,2}$ Eq. (10) has the form
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (16)
-Taking into account the boundary conditions (15) we obtain
-$E\sim\left(V-V_{1,2}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (17)
-If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
-$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (18)
-Eq. (10) can be specified to
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (19)
-For brevity we’ll call Eq. (18) and (19) the cubic case. Not that in this case
-$\frac{1}{LC_{0}u^{2}(V)}=\alpha+2\beta V+3\gamma V^{2}.$ (20)
+Also, from (10a) and (14) follows that in the vicinity of $V_{1,2}$
+$E^{2}\sim\left(V-V_{1,2}\right)^{2}.$ (17)
+The conditions (16) and (17) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
 II.2 The kinks
-With four boundary conditions on $E(V)$ supplied by (15) and (17) for the kink problem of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two being connected with it.
-One such connection we could obtain at much earlier stage.
-In fact, integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
-$\frac{1}{U^{2}L}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (21)
+With four boundary conditions specifying both the position and the velocity of the particle at $V=V_{1}$ and $V=V_{2}$, of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two are connected with it.
+One such constraint we could obtain at much earlier stage.
+By integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
+$\frac{1}{U^{2}}=L\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (18)
 where $Q_{i}=Q(V_{i})$.
 If $Q(V)$ is a polynomial function of $V$, from (17) follows that the kink can be presented as
-$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2}.$ (22)
-In the cubic case, calculating the integral in (21) we obtain
-$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right).$ (23)
+$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2},$ (19)
+where $P_{1}(V)$ is some polynomial function.
+If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
+$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right)$ (20)
+(for brevity we’ll call Eq. (20) the cubic case), then
+$\frac{1}{u^{2}(V)}$ $=LC_{0}\left(\alpha+2\beta V+3\gamma V^{2}\right),$ (21a)
+$\frac{1}{U^{2}}$ $=LC_{0}\left[\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right)\right].$ (21b)
 We also can specify (19) to
-$E^{2}(V)=\psi^{2}(V-V_{1})^{2}(V-V_{2})^{2},$ (24)
-where $\psi$ is a constant. Substituting (24) into (19) and using Vieta’s formula we obtain after simple algebra
+$E(V)=\psi(V-V_{1})(V-V_{2}),$ (22)
+where $\psi$ is a constant (positive, if $V_{1}>V_{2}$, and negative in the opposite case). Substituting (22) into (13) and using Vieta’s formula we obtain after simple algebra
 $\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (23a)
 $V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}.$ (23b)
-To find $V(Z)$ we have to solve equation
-$\frac{dV}{dZ}=E(V).$ (26)
-The solution is
-$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (27)
-After finding $V(Z)$ we can from (2) obtain the current
-$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2UL}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (28)
+Thus the necessary condition for the existence of the kink is the inequality $\gamma>0$, that is the speed of sound $u^{2}(V)$ should be a convex function of $V$.
+To find $V(Z)$ we have to solve Eq. (9). The solution is
+$V(Z)=\frac{V_{1}+V_{2}}{2}-\frac{V_{1}-V_{2}}{2}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (24)
+After finding $V(Z)$ we can from (1b) obtain the current
+$I(Z)=-\frac{\left(V_{1}-V_{2}\right)}{2UL}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (25)
 II.3 The solitons
@@
 Because $V_{3}$ is the extremal value of $V$ we have
-$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0\Rightarrow E(V_{3})=0.$ (29)
-Hence
-$E^{2}\sim\left(V-V_{3}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{3}.$ (30)
+$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0,\hskip 28.45274pt\left(\frac{d^{2}V}{dZ^{2}}\right)_{V=V_{3}}\neq 0.$ (26)
+Hence from (9) and (10a) follows that in the vicinity of $V_{3}$
+$E^{2}\sim\left(V-V_{3}\right).$ (27)
 Recalling our mechanical analogy we see that the soliton corresponds to the motion of the fictitious particle with zero initial coordinate and velocity at $V=V_{1}$ and zero coordinate (but not velocity) at $V=V_{3}$.
@@
 This is why $E(V)$ for the soliton will later turn out to be a double-valued function.
-With three boundary conditions on $E(V)$ supplied by (15) and (29) for the soliton problem of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters.
+With the position and the velocity of the particle at $V=V_{1}$ and its position at $V=V_{3}$ being specified, of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters.
 Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
@@
 $E^{2}(V)=P_{2}(V)(V-V_{1})^{2}(V-V_{3}),$ (29)
-where $P_{1}(V)$ and $P_{2}(V)$ are some polynomial functions.
+where $P_{2}(V)$ is some polynomial function.
 In the cubic case, calculating the integral in (28) we obtain
-$\frac{1}{LC_{0}U^{2}}=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right).$
-We also can specify (32) (generalize (24)) to
+$\frac{1}{U^{2}}=LC_{0}\left[\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right)\right].$
+We also can specify (29) to
 $E^{2}(V)=\psi(V-V_{1})^{2}(V-V_{3})(V-V_{4}),$ (31)
 where $\psi$ is a constant. Substituting (31) into (13) and using Vieta’s formula we obtain after simple algebra
-$\psi^{2}=6U^{2}LC_{0}\gamma,$ (35a)
+$\psi=6U^{2}LC_{0}\gamma,$ (32a)
 $2V_{1}+V_{3}+V_{4}=-\frac{4\beta}{3\gamma}.$ (32b)
@@
 Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (20).
-In this case, excluding $V_{4}$, we obtain that (34) and (II.3) are modified to
+In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (31) and (II.3) are modified to
 $E^{2}(V)$ $=\phi(V-V_{1})^{2}(V-V_{3}),$ (33a)
-$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta,$ (36b)
-$\frac{1}{LC_{0}U^{2}}$ $=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right).$ (36c)
-From (II.3) we see that the soliton exists both for positive and negative $\beta$.
-For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
-From (36c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation (20) with $\gamma=0$, in line with our previous general statement that the soliton is supersonic.
-To find $V(Z)$ we have to solve equation
-$\frac{dV}{dZ}=E(V).$ (37)
-The solution of (37) is
+$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta.$ (33b)
+From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
+To find $V(Z)$ we have to solve Eq. (9). The solution is
 $V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (34)
-Similar to what was done in the previous subsection, we can use (2) to find the current and obtain
-$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (39)
+After finding $V(Z)$ we can from (1b) obtain the current
+$I(Z)=\frac{V_{3}-V_{1}}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (35)
 II.4 The quasi-continuum approximation
@@
 Figure 2: Discrete transmission line with shunted inductors.
-In this case (1c) is modified to
-$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}+\frac{1}{R}\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (42)
-hence Eq. (3) takes the form
-$L\frac{\partial I^{c}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}-\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial Z^{2}}.$ (43)
-(we assume that both the higher order derivatives with respect to $Z$ and $1/R$ are small, so the discrete second order derivative with respect to $Z$ in the r.h.s.
-of (42) we presented as the continuous second order derivative).
-From (1a) and (43) we can obtain Eq. (4) modified due to the presence of ohmic resistors
-$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (44)
-For the travelling wave, combining (1a) and (43), we obtain after integration closed equation for $V(Z)$
-$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=U^{2}LQ-V+F,$ (45)
-where $F$ is the constant of integration.
+In this case the circuit equations are modified to
+$L\frac{dI_{n,n+1}}{dt}$ $=\left(1+R_{C}C\frac{d}{dt}\right)\left(V_{n}-V_{n+1}\right),$ (38a)
+$I^{c}_{n}$ $=\left(1+\frac{L}{R_{L}}\frac{d}{dt}\right)\left(I_{n-1,n}-I_{n,n+1}\right)$ (38b)
+(Eq. (1a) remains as it was). Hence (3) is modified to
+$L\frac{d^{2}Q_{n}}{dt^{2}}=\left(1+\frac{L}{R_{L}}\frac{d}{dt}\right)\left(1+R_{C}C\frac{d}{dt}\right)$
+$\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (39)
+and (4) is modified due to the presence of ohmic resistors to
+$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\tau_{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}},$ (40)
+where
+$\tau_{R}=\frac{L}{R_{L}}+R_{C}C.$ (41)
+In deriving (40) we assumed that both the influence of the higher order derivatives with respect to $Z$ and the dissipation are small, so we ignored the terms which which correspond to the products of these effects.
+For the travelling wave we obtain from (40) after double integration the generalization of (8)
+$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=U^{2}LQ(V)-V+F.$ (42)
 Equation (42) is Newton equation, describing motion with friction of the fictitious particle ($V$ being the coordinate of the particle and $Z$ the ”time”) in the force field given by the r.h.s.
@@
 It will be convenient to present Eq. (42) as
-$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (46)
+$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (43)
 where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
-$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (47)
+$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q(V)+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (44)
 The constraint for the travelling wave asymptotic boundary conditions is the inequality
@@
 […]
-When both $\beta>0$ and $\gamma>0$, the condition (49) means just $|V_{1}|>|V_{2}|$.
+For $\gamma>0,\beta\geq 0$, the condition (46) means just $|V_{1}|>|V_{2}|$.
 The point $V=V_{1}$ should be the position of unstable equilibrium.
 The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium.
-We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ).
-The stable equilibrium waves we called shocks, and the unstable equilibrium - kinks.
-There are several important distinctions between these two types of waves.
-Unstable equilibrium at both at the point $V=V_{1}$ and $V=V_{2}$ means that differentiating the r.h.s.
-of (45) we recover equation (12).
-The kink, similar to what was in the lossless case, is supersonic both with respect to the region before the kink and the region behind the kink.
-Since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation.
-Stable equilibrium at the point $V=V_{2}$ means
-$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{2}}-1<0.$ (50)
-Inequalities (50) and (12) (the latter only for $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
-To clarify the distinction between the shocks and the kinks let us bring together the formulas for the travelling wave speed and the sound speed for the cubic case
-$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right),$ (51a)
-$\frac{1}{LC_{0}u^{2}(V_{1,2})}=\alpha+2\beta V_{1,2}+3\gamma V_{1,2}^{2}.$ (51b)
-We realize that for $\beta>0$ and $\gamma>0$, for $V_{1}$ and $V_{2}$ having the same sign the travelling wave is always the shock.
-For $V_{1}$ and $V_{2}$ having opposite signs, we can have both the kink and the shock, depending upon the parameters.
-To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (10)
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L\frac{dQ(V)}{dV}-1.$ (52)
-The boundary conditions (11) lead to the boundary conditions (15) for $E(V)$, like it was in the lossless case. Equation (52) in the vicinity of $V_{1,2}$ has the form
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=\text{const},$ (53)
-which contains additional term in the l.h.s. with respect to (16), but we still obtain the condition (17). We can also repeat verbatim the part of Section II.2 leading to Eqs. (21) and (22).
-In the cubic case Eq. (52) becomes
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (54)
-We regain Eq. (23), and there exists kogan4 a simple polynomial solution of (54) in the form (24), which we repeat here
-$E(V)=\psi(V-V_{1})(V-V_{2}),$ (55)
-where $\psi$ is a constant (positive for positive $V_{1}$).
-Substituting (55) into (54) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
-$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (56a)
-$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4UL}{\psi R}.$ (56b)
-Note that Eqs.
-(25a), (35a) and (56a) are all identical, hence all the waves we consider in this paper can exist only if $\gamma>0$.
-Note also that by integrating (54) we can obtain Abel equation of the second kind, and (56b) is the condition of the solvability of the equation polyanin .
-Equation (55) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (27)).
-And the the shocks are of the same shape also.
-However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq.
-(56b) instead of (25b)) and so does the kink (the shock) speed, in spite of the validity in all the cases of Eq.
-(23).
-The boundary in the phase space between the kinks and the shocks can be easily obtained from (III) in the case when $\beta=0$. Equating the speeds we obtain that this boundary corresponds to
+(We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ).
+The waves corresponding to the stable equilibrium at the point $V=V_{2}$ we called the shocks, to the unstable equilibrium – the kinks.
+Differentiating (44) we obtain the r.h.s. of (49) (see below)
+$-\frac{d^{2}\Pi(V)}{dV^{2}}=\frac{U^{2}}{u^{2}(V)}-1;$ (47)
+the point $V=V_{i}$ being the point of the stable/unstable equilibrium corresponds to the travelling wave being subsonic/supersonic with respect to the background.
+Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{2}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
+Let us return to (II.2) giving the sound and the kink/shock speeds for the cubic case.
+The kinks/shocks phase diagram in the space $V_{1},V_{2}$ can be obtained by comparing $U^{2}$ and $u^{2}(V_{2})$.
+Consider the case $\gamma>0,\beta\geq 0$.
+If $V_{1}$ and $V_{2}$ have the same sign, the travelling wave is always the shock.
+In the opposite case, the wave can be either the kink or the shock, depending upon $V_{1}$ and $V_{2}$.
+The boundary in the phase space between the kinks and the shocks is especially simple for $\beta=0$
 $V_{1}+2V_{2}=0$ (48)
 (very much similar to what we had for the Josephson transmission line kogan4 ).
+The most important distinction of the dissipative kinks from the shocks: since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation (see below).
+The shocks don’t need such find tuning – they are generic.
+To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (11)
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=\frac{U^{2}}{u^{2}(V)}-1.$ (49)
+At this stage we can repeat verbatim the part of Section II.2 leading to (22), hence the dissipative kinks and the shocks are of the same shape (given by (24)) as the kinks in the lossless line.
+However, Eq. (II.2) is generalized to
+$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (50a)
+$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (50b)
+hence the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq.
+(50b) instead of (23b)) and so does the kink (the shock) speed and amplitude, in spite of the validity in all the cases of Eq.
+(21b).
 We need to emphasise the different meaning of the condition (50b) for the kinks and the shocks.
-For the kinks, (56b) can be understood as the constraint on $V_{2}$ (for a given value of $V_{1}$ and ohmic resistance $R$).
+For the kinks, (50b) is the constraint on $V_{1},V_{2},R$.
 And it happens that the kink satisfying such constraint is described by the elementary function.
 For the shock, (50b) is just the condition for it being described by an elementary function.
-But, of course, the shock with any other value of $V_{2}$ (satisfying (48) also exists.
+But, of course, the shock with any other values of $V_{1},V_{2},R$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (45)).
+Note also that by integrating (49) we can obtain Abel equation of the second kind, and (50b) is the condition of the solvability of the equation polyanin .
 Let us assume that $V_{1}$ and $V_{2}$ are so close to each other, that we can keep in the r.h.s. of (20) only the first two terms, thus simplifying (49) to
-$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1.$ (58)
-The solution of (58) satisfying the boundary conditions (15) and (29) is
-$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right],$ (59)
-where $\Delta V\equiv V_{1}-V_{2}$ (we assumed for the sake of definiteness that $V_{1},V_{2}>0$). Substituting (59) into (58) we obtain
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1$ (51)
+(we assume for the sake of definiteness that $\beta>0$ and $V_{1},V_{2}>0$). The elementary solution of (51) satisfying the boundary conditions (16) is of the form kogan4
+$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(V_{1}-V_{2}\right)^{1/2}\right].$ (52)
+Substituting (52) into (51) we obtain
 $\chi^{2}$ $=8U^{2}LC_{0}\beta$ (53a)
-$\frac{UL}{R}$ $=\frac{5\chi}{24}\left(\Delta V\right)^{1/2}.$ (60b)
-The function $E(V)$ being found, we can find the function $V(Z)$ by solving equation
-$\frac{dV}{dX}=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right].$ (61)
-The solution is
-$V(Z)=V_{2}+\frac{\Delta V}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (62)
-According to what was said after Eq.
-(III), Eq.
-(62) describes the shock.
-And we have to repeat here what was said after Eq.
-(III).
-Equation (60b) is just the condition for the shock to be expressed by an elementary function.
-The shock exists for any other value of $V_{2}$.
+$U\tau_{R}$ $=\frac{5\chi}{24}\left(V_{1}-V_{2}\right)^{1/2}.$ (53b)
+Similar to Eq. (50b) in the case of shocks, Eq. (53b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $\Delta V$.
+The function $E(V)$ being found, we can find the function $V(Z)$ by solving Eq. (9). The solution is
+$V_{2}(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (54)
 IV Conclusions
@@
 Acknowledgements.
-I am grateful to B.
-Malomed for the discussion.
+I am grateful to A.
+M.
+Kamchatnov, B.
+Malomed and A.
+Vainchtein for the discussion.
 The paper was finalized during the Author’ participation in Workshop on Twistronics and Moire Materials: Bridging Theory and Experiments, ICTP, Trieste.

### v6

@@
 Abstract
-We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors.
-The travelling waves with finite support in such systems are the kinks and the solitons, which speeds and profiles were calculated.
-We also studied lossy transmission lines.
-The travelling waves with finite support in such systems are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
+We studied discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors (and possibly resistors).
+The localised travelling waves in the lossless transmission lines are the kinks and the solitons, which speeds and profiles were calculated.
+The localised travelling waves in the lossy transmission lines are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
 I Introduction
@@
 The discrete transmission line is constructed from the identical inductors and nonlinear capacitors is shown on Fig. 1.
-Figure 1: Discrete transmission line.
+Figure 1: The lossless discrete transmission line.
 We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
@@
 of the equation.
-We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
+We’ll consider localised travelling waves, that is impose upon $V(Z)$ the boundary conditions
 $\lim_{Z\to-\infty}V(Z)=V_{1},\hskip 28.45274pt\lim_{Z\to+\infty}V(Z)=V_{2}.$ (14)
@@
 III The dissipative kinks and shocks
-Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
-Figure 2: Discrete transmission line with shunted inductors.
+Let us ask ourselves what happens with the kinks when the transmission line becomes lossy?
+To answer this question let us include in the system the resistors $R_{L}$ shunting the inductors and the resistors $R_{C}$ in series with the nonlinear capacitors (see Fig.
+2).
+Figure 2: The lossy discrete transmission line.
 In this case the circuit equations are modified to
@@
 The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium.
-(We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ).
+(We met with similar dichotomy while studying localised travelling waves in Josephson transmission line kogan4 ).
 The waves corresponding to the stable equilibrium at the point $V=V_{2}$ we called the shocks, to the unstable equilibrium – the kinks.
@@
 IV Conclusions
-We considered finite support travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors.
+We considered localised travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors.
 We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line.
 We also studied lossy transmission lines.
-The travelling waves with finite support in such systems are the dissipative kinks and the shocks.
+The localised travelling waves in such systems are the dissipative kinks and the shocks.
 We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.

### v7

@@
-On the kinks, the solitons and the shocks in discrete nonlinear transmission line
+Exact analytical solutions for the kinks, the solitons and the shocks in discrete nonlinear transmission line with nonlinear capacitance
 Abstract
 We studied discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors (and possibly resistors).
-The localised travelling waves in the lossless transmission lines are the kinks and the solitons, which speeds and profiles were calculated.
-The localised travelling waves in the lossy transmission lines are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
+The localised travelling waves in the lossless transmission lines are the kinks and the solitons, the localised travelling waves in the lossy transmission lines are the dissipative kinks and the shocks.
+The speeds and profiles of all these waves were calculated.
 I Introduction
@@
 Figure 1: The lossless discrete transmission line.
-We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
+We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n-1,n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
 $\frac{dQ_{n}}{dt}$ $=I^{c}_{n},$ (1a)
@@
 In this paper we are interested in the travelling wave solutions, for which the dependence of all the variables upon $Z$ and $t$ is of the form
-$X(Z,t)=X(Z-Ut),$ (5)
+$X(Z,t)=X(Z+Ut),$ (5)
 where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This from (4) we obtain an ordinary differential equation
 $U^{2}L\frac{d^{2}Q(V)}{dZ^{2}}=\frac{d^{2}V}{dZ^{2}}+\frac{1}{12}\frac{d^{4}V}{dZ^{4}}.$ (6)
-Performing an integration we get
-$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
-One way to solve Eq. (7) is to perform the second integration to obtain
+After finding $V(Z)$ we can from (1b) obtain the current
+$I(Z)=-\frac{V(Z)}{UL}+\text{const}.$ (7)
+Performing an integration in (6) we can get
+$\frac{1}{12}\frac{d^{3}V}{dZ^{3}}=U^{2}L\frac{dQ}{dZ}-\frac{dV}{dZ}.$ (8)
+Notice that equation with the same form as (8) one would obtain after making the travelling wave ansatz in the generalised Korteweg-de Vries-modified Korteweg-de Vries (KDV-mKDV) equation.
+We may perform the second integration to obtain
 $\frac{1}{12}\frac{d^{2}V}{dZ^{2}}=U^{2}LQ(V)-V+F,$ (9)
-where $F$ is the constant of integration.
-However, in this Section we’ll use another method of solution of Eq. (7). After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, we are motivated to introduce the new dependent variable
-$E=\frac{dV}{dZ}.$ (9)
-Considering $V$ as the independent variable, taking into account that
-$\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (10a)
-$\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (10b)
-and dividing both parts by the common multiplier $E$ in both parts, we may reduce the order of (7) and write it down as
+where $F$ is the constant of integration. Introducing the new dependent variable
+$E=\frac{dV}{dZ},$ (10)
+considering $V$ as the independent variable and taking into account that
+$\frac{d^{2}V}{dZ^{2}}=\frac{1}{2}\frac{dE^{2}}{dV},$ (11)
+we may present (9) as
+$\frac{1}{24}\frac{dE^{2}}{dV}=U^{2}LQ(V)-V+F.$ (12)
+However, for technical reasons, it will be more convenient to return to (8) and avoid the integration of the equation. After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, taking into account that
+$\frac{d^{3}V}{dZ^{3}}=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}},$ (13)
+and dividing both parts of (8) by the common multiplier $E$ in both parts, we may reduce the order of the equation and write it down as
 $\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\frac{dQ(V)}{dV}-1.$ (14)
@@
 $E\left(V_{1,2}\right)=0.$ (19)
-Also, from (10a) and (14) follows that in the vicinity of $V_{1,2}$
+Also, from (11) and (17) follows that in the vicinity of $V_{1,2}$
 $E^{2}\sim\left(V-V_{1,2}\right)^{2}.$ (20)
 The conditions (19) and (20) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
+The main result of the Section II.1 is Eq.
+(14).
+The natural question arises: why we were not happy with Eq.
+(9) (which can be easily integrated in quadratures) and introduced instead of it the former equation.
+The detailed answer to this question will be given in Section III, but we can formulate it announce it right now.
+The main aim of the present paper is obtaining exact analytical solutions for the kinks and the shocks in the lossy transmission line.
+When the loss terms are included into the equations, (14) can be integrated analytically much easier than (9).
 II.2 The kinks
@@
 $V(Z)=\frac{V_{1}+V_{2}}{2}-\frac{V_{1}-V_{2}}{2}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (27)
-After finding $V(Z)$ we can from (1b) obtain the current
-$I(Z)=-\frac{\left(V_{1}-V_{2}\right)}{2UL}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (25)
 II.3 The solitons
@@
 $\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0,\hskip 28.45274pt\left(\frac{d^{2}V}{dZ^{2}}\right)_{V=V_{3}}\neq 0.$ (28)
-Hence from (9) and (10a) follows that in the vicinity of $V_{3}$
+Hence from (10) and (11) follows that in the vicinity of $V_{3}$
 $E^{2}\sim\left(V-V_{3}\right).$ (29)
@@
 Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (23).
-In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (31) and (II.3) are modified to
+In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (33) and (34) are modified to
 $E^{2}(V)$ $=\phi(V-V_{1})^{2}(V-V_{3}),$ (35a)
 $\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta.$ (35b)
-From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
+From (35) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
 To find $V(Z)$ we have to solve Eq. (10). The solution is
 $V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (36)
-After finding $V(Z)$ we can from (1b) obtain the current
-$I(Z)=\frac{V_{3}-V_{1}}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (35)
-II.4 The quasi-continuum approximation
+II.4 How good is the quasi-continuum approximation?
 We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper.
@@
 $U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (38)
-Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
+Looking at the results of the previous Sections we understand that the fifth order derivative in Eq.
+(38) can be discarded if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
 The small numerical coefficient before the sixth order derivative in the r.h.s.
 of (38) means that the small parameters shouldn’t be actually very small.
-With all that, one can expect that Eq.
-(37) will allow to study the cases of stronger nonlinearity than Eq.
-(7).
+With all that we have to admit that the transition from the discrete to the continuous system is an uncontrollable approximation.
+More specifically, the mere fact of existence of the travelling wave of the form (5) in the discrete system should be proved analytically or, at least, numerically.
+We hope to return to this issue in the future.
 III The dissipative kinks and shocks
@@
 where
-$\tau_{R}=\frac{L}{R_{L}}+R_{C}C.$ (41)
-In deriving (40) we assumed that both the influence of the higher order derivatives with respect to $Z$ and the dissipation are small, so we ignored the terms which which correspond to the products of these effects.
-For the travelling wave we obtain from (40) after double integration the generalization of (8)
+$\tau_{R}\equiv\frac{L}{R_{L}}+R_{C}C.$ (42)
+In deriving (41) we assumed that the dissipation is small, so we ignored the terms which correspond to the product of the dissipation coefficients and the higher order derivatives with respect to $Z$.
+For the travelling wave we obtain from (41) after integration the generalization of (8)
+$\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+U\tau_{R}\frac{d^{2}V}{dZ^{2}}=U^{2}L\frac{dQ}{dZ}-\frac{dV}{dZ}.$ (43)
+Similar to what we said in Section II.1, notice that equation with the same form as (43) one would obtain after making the travelling wave ansatz in the generalised Korteweg-de Vries-modified Korteweg-de Vries (KDV-mKDV) equation with dissipation.
+Performing the second integration, we obtain the generalization of (9)
 $\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=U^{2}LQ(V)-V+F.$ (44)
+From (39) we obtain that in the present case the sum of the currents flowing through the nonlinear inductance and the ohmic resistance $R_{L}$ is given by the equation
+$I(Z)=-\frac{V(Z)}{UL}-\frac{\tau_{R}}{L}\frac{dV(Z)}{dZ}+\text{const}.$ (45)
 Equation (44) is Newton equation, describing motion with friction of the fictitious particle ($V$ being the coordinate of the particle and $Z$ the ”time”) in the force field given by the r.h.s.
@@
 $\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (46)
-where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
+where
 $\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q(V)+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (47)
@@
 the point $V=V_{i}$ being the point of the stable/unstable equilibrium corresponds to the travelling wave being subsonic/supersonic with respect to the background.
-Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{2}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
-Let us return to (II.2) giving the sound and the kink/shock speeds for the cubic case.
+Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
+Let us return to (24) giving the sound and the kink/shock speeds for the cubic case.
 The kinks/shocks phase diagram in the space $V_{1},V_{2}$ can be obtained by comparing $U^{2}$ and $u^{2}(V_{2})$.
@@
 The shocks don’t need such find tuning – they are generic.
-To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (11)
+To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there starting from Eq. (43), we obtain modified Eq. (14)
 $\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=\frac{U^{2}}{u^{2}(V)}-1.$ (52)
-At this stage we can repeat verbatim the part of Section II.2 leading to (22), hence the dissipative kinks and the shocks are of the same shape (given by (24)) as the kinks in the lossless line.
-However, Eq. (II.2) is generalized to
-$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (50a)
-$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (50b)
-hence the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq.
-(50b) instead of (23b)) and so does the kink (the shock) speed and amplitude, in spite of the validity in all the cases of Eq.
-(21b).
-We need to emphasise the different meaning of the condition (50b) for the kinks and the shocks.
-For the kinks, (50b) is the constraint on $V_{1},V_{2},R$.
-And it happens that the kink satisfying such constraint is described by the elementary function.
-For the shock, (50b) is just the condition for it being described by an elementary function.
-But, of course, the shock with any other values of $V_{1},V_{2},R$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (45)).
-Note also that by integrating (49) we can obtain Abel equation of the second kind, and (50b) is the condition of the solvability of the equation polyanin .
+Considering again the cubic case we can easily check up that (25) is still the solution of Eq. (52), with $\psi$ being still given by Eq. (26a), but Eq. (26b) is generalized to
+$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (53)
+Hence the dissipative kinks and the shocks are of the same shape (given by (27)) as the kinks in the lossless line.
+Now we understand what all the trouble of introducing the new dependent and independent variables was for.
+We were able to integrate analytically Eq.
+(52) (while we didn’t see any way to integrate analytically Eq.
+(44)).
+We need to emphasise the different meaning of the condition (53) for the kinks and the shocks.
+For the kinks, (53) is the condition of the solution existence (and the solution happens to be described by the elementary function).
+For the shock, (53) is just the condition for the solution to be described by an elementary function.
+But, of course, the shock with any other values of $V_{1},V_{2},\tau_{R}$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (48)).
+So the obtained shock can be called the particular shock of the first kind.
+Note also that by integrating (52) we can obtain Abel equation of the second kind, and (53) is the condition of the solvability of the equation polyanin .
 Let us assume that $V_{1}$ and $V_{2}$ are so close to each other, that we can keep in the r.h.s. of (23) only the first two terms, thus simplifying (52) to
@@
 $U\tau_{R}$ $=\frac{5\chi}{24}\left(V_{1}-V_{2}\right)^{1/2}.$ (56b)
-Similar to Eq. (50b) in the case of shocks, Eq. (53b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $\Delta V$.
+Similar to Eq.
+(53) in the case of shocks, Eq.
+(56b) is just the condition for the shock to be expressed by an elementary function.
+The shock exists for any other value of $V_{1}-V_{2}$.
+So the obtained shock can be called the particular shock of the second kind.
 The function $E(V)$ being found, we can find the function $V(Z)$ by solving Eq. (10). The solution is
-$V_{2}(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (54)
+$V(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (57)
 IV Conclusions
@@
 The localised travelling waves in such systems are the dissipative kinks and the shocks.
-We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.
+The velocity of the kink, both in the lossless and in the lossy line, and of the shock in the lossy line is given by the same equations
+$\frac{1}{U^{2}}=L\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}},$ (58)
+where $Q_{1},V_{1}$ and $Q_{2},V_{2}$ are the asymptotic values of the capacitor charge and the voltage on both ends of the transmission line.
+To our opinion the most important equation of the paper is Eq. (52) describing the lossy transmission line
+$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=\frac{U^{2}}{u^{2}(V)}-1.$ (59)
+When the charge of the non-linear capacitor and the voltage on it are connected by the equation
+$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (60)
+the profiles of the kink, both in the lossless and in the lossy line, and of the particular shock of the first kind in the lossy line, can be easily calculated on the basis of the solution of Eq. (59)
+$E(V)=\psi(V-V_{1})(V-V_{2}),$ (61)
+and are given by the same equation
+$V(Z)=\frac{V_{1}+V_{2}}{2}-\frac{V_{1}-V_{2}}{2}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right],$ (62)
+where $\psi^{2}=6U^{2}LC_{0}\gamma$, and the condition on the asymptotic values of the voltage on both ends of the transmission line $V_{1}$ and $V_{2}$ are given by Eq.
+(53).
+The profile of the particular shock of the second kind is given by Eq.
+(57.
 Acknowledgements.